The notation ΣXY refers to ______.
1. (12+33+44+46+52+60+75) x (11+39+48+50+54+63+70)
2. (ΣX)(ΣY)
3. (12x11)+(33x39)+(44x48)+(46x50)+(52x54)+(60x63)+(75x70)
4. (12+11) (33+39) (44+48) (46+50) (52+54) (60+63) (75+70)
1
Expert's answer
2020-08-26T17:06:38-0400

"\sum XY = (x_1 \cdot y_1) + (x_2 \cdot y_2) + ...+ (x_n \cdot y_n)"

The line which follows this pattern is 3.

Answer: 3. (12x11)+(33x39)+(44x48)+(46x50)+(52x54)+(60x63)+(75x70)